OpenStudy (anonymous):

if "a" and "b" are acute angles with {sin a = 12/13} and {sin b= 8/17}, what is the value of tan (a +b)?

OpenStudy (dumbcow):

tan(a+b) = (tan a + tan b)/ (1-tan(a)tan(b)) sin = opp/hyp find adjacent side using pythagorean thm 13^2 - 12^2 = 5^2 17^2 - 8^2 = 15^2 tan = opp/adj tan a = 12/5 tan b = 8/15 sub into formula for tan(a+b)

OpenStudy (anonymous):

ok ty i got it now.

OpenStudy (dumbcow):

yw :)